Calculation of Bohr's atomic model of energy of electron in nth orbit.

1
Expert's answer
2021-11-30T09:50:07-0500

The energy of electron in "n"th orbit can be calculated with the following equation:


"E_n=-\frac{2\pi^2me^4Z^2}{n^2h^2(4\pi\epsilon_0)^2}=-13.6\frac{Z^2}{n^2}\space\text{ eV}."